OpenStudy (anonymous):

6. The Martians ask you to explain one last thing, Ultimate Math Ambassador. They ask you to create a new function, h(x). Then assign any number to x. Using complete sentences, explain whether f(h(x)) and h(f(x)) will always result in the same number. You will use the function f(x) that you created in problem number 2. This is my h function h(x)=10x-5 my f(x) f(x)=9x+6

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

i need help

OpenStudy (anonymous):

1. Two Martians, Splott and Fizzle, have solved the equation 2x + 4 = –3x + 14. Examine the work of Splott and Fizzle. Identify any errors in the Martians’  calculations and explain, using complete sentences, what corrections they should make. One mistake that Splott made was that he should have subtracted 4 not added it to +4, done the same thing to 14, subtract 4 not add it then he should change 18 to 10 and change the ange to x=2 since 10/5 is 2 For Fizzle he just did te math wrong for 4-14 he put -12, he should have put -10. which would then make the problem -10=-5x. Then he should divide -5by both sides to get the answer 2. Splott Fizzle 2x + 4 = –3x + 14 2x + 4 = –3x + 14 2x + 4 + 3x = –3x + 14 + 3x 2x + 4 – 2x = –3x + 14 – 2x 5x + 4 = 14 4 = 14 – 5x 5x + 4 + 4 = 14 + 4 4 – 14 = 14 – 5x – 14 5x = 18 –12 = –5x 5x/5=18/5 –12 + 5 = –5x + 5 X= 18/5 –7 = x 2. Create your own function to teach the Martians about functions. Your function must contain at least two different operations. f(x)=6x+9 3. Using complete sentences, prove to Splott and Fizzle that your function is a legitimate function. The reason this is and legit function is that for every x value there is a unique y value. 4. Using your function, explain to the Martians how to solve for f(3). Show your work and explain each step using complete sentences. F(3)=6(3)+9 f(3)=18+9 f(3)=27 5. Using complete sentences, describe to the Martians how to find the inverse of your function. First you would take the y value 27 and replace f(3) with it. And then make the equation 27=6x+9. Then you subtract 9 from both sides, like so 27-9=6x+9-9. You are then left with this, 27-9=6x. Then you divide 6 by both sides, to get this, 27-9/6=6x/6. Then after tyou do the division and you are left with 27-9/6=x. 6. The Martians ask you to explain one last thing, Ultimate Math Ambassador. They ask you to create a new function, h(x). Then assign any number to x. Using complete sentences, explain whether f(h(x)) and h(f(x)) will always result in the same number. You will use the function f(x) that you created in problem number 2. h(x)=10x+5

OpenStudy (anonymous):

i really only need help with the last question

OpenStudy (anonymous):

i dont understand how to start it

ganeshie8 (ganeshie8):

You have : ``` h(x)=10x-5 f(x)=9x+6 ```

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

you wanto to find whether ` f(h(x))` and `h(f(x)) ` result in the same number or not for some value of `x`. Is that right ?

OpenStudy (anonymous):

yes and i dont understand how to

OpenStudy (anonymous):

or how to explain it

ganeshie8 (ganeshie8):

Lets find `f(h(x))` first, we need to pick some value for `x`, lets say x=1 ?

ganeshie8 (ganeshie8):

``` h(x)=10x-5 f(x)=9x+6 ``` to find `h( f(1) ) ` , you need to work `f(1)` first

ganeshie8 (ganeshie8):

f(1) = ?

OpenStudy (anonymous):

f(1)=9(1)+6

ganeshie8 (ganeshie8):

yes, simplify

OpenStudy (anonymous):

f(1)=9+6 -6 f(1)=9

ganeshie8 (ganeshie8):

NO. f(1)=9(1)+6 = 9 + 6 = 15 so the value of f(1) is 15

OpenStudy (anonymous):

oh ok so f(1)=15

ganeshie8 (ganeshie8):

So, h( f(1) ) = h(15) = ?

ganeshie8 (ganeshie8):

whats the value of h(15) ?

OpenStudy (anonymous):

see this is what confuses me

ganeshie8 (ganeshie8):

you have ``` h(x)=10x-5 ``` plugin x = 15 to get h(15) : ``` h(15) = 10*15 - 5 = ? ```

OpenStudy (anonymous):

h(15)=10*15-5 =150-5 =145 h(15)=145

ganeshie8 (ganeshie8):

Yes! so overall we have : h( f(1) ) = h( 15 ) = 145 fine ?

ganeshie8 (ganeshie8):

similarly, can you find the value of `f( h(1) )` ?

OpenStudy (anonymous):

f(h(1))=f(145) or am i wrong

OpenStudy (anonymous):

or f(h(1))=f(15) Which one is right?

ganeshie8 (ganeshie8):

to find `f( (h(1) )`, you need to work `h(1)` first h(1) = ?

OpenStudy (anonymous):

h(1)=10(1)-5 =10-5 h(1)=5

ganeshie8 (ganeshie8):

yes! f( h(1) ) = f( 5 ) = ?

ganeshie8 (ganeshie8):

whats the value of f(5) ?

OpenStudy (anonymous):

f(5)=10(5)-5 =50-5 f(5)=45

ganeshie8 (ganeshie8):

Correct! so f( h(1) ) = f( 5 ) = 45

ganeshie8 (ganeshie8):

we have got : `h(f(1)) = 145` and `f(h(1)) = 45` Clearly they're NOT same right ?

OpenStudy (anonymous):

True. but now i have to fiigure out how to put that all into complete sentences

ganeshie8 (ganeshie8):

show the work above and conclude that `h(f(x))` and `f(h(x))` need not be same

OpenStudy (anonymous):

oh ok thanks man

ganeshie8 (ganeshie8):

yw :)